OpenStudy (anonymous):

what is the sum of an infinite geometric series with a term of 6 and a common ratio of 1/2

sam (.sam.):

1st term of 6?

sam (.sam.):

\[S_\infty =\frac{a_1}{1-r}\] \[S_\infty= \frac{6}{1-\frac{1}{2}}=?\]

OpenStudy (anonymous):

thanks!

sam (.sam.):

welcome :)

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

HEY HEY !!!!!!!!!!!!!!!!!!!!!!

sam (.sam.):

yes?